OpenStudy (anonymous):

what is x+x ... is it x^2

OpenStudy (anonymous):

No its 2x.

OpenStudy (anonymous):

x times x is x^2

OpenStudy (anonymous):

he's right

OpenStudy (anonymous):

oooh thank you :)

OpenStudy (anonymous):

x*x = x^2 x + x = 2x

OpenStudy (asnaseer):

it may help you to tink of "x" as representing some real life object - say books. so: x + x would mean 1 book plus another book, which makes 2 books.

OpenStudy (asnaseer):

**think

OpenStudy (asnaseer):

 where as if you had three rows of books, with 3 books in each row, then this would represent: x * x = x^2 which would be 3 times 3 = 3 squared = 9 books